OpenStudy (anonymous):

Find the center, vertices, and foci of the ellipse with equation. x squared divided by 81 plus y squared divided by 225 equals 1 Center: (0, 0); Vertices: (-15, 0), (15, 0); Foci: (0, -9), (0, 9) Center: (0, 0); Vertices: (0, -15), (0, 15); Foci: (-9, 0), (9, 0) Center: (0, 0); Vertices: (0, -15), (0, 15); Foci: (0, -12), (0, 12) Center: (0, 0); Vertices: (-15, 0), (15, 0); Foci: (-12, 0), (12, 0)

Nnesha (nnesha):

\[\huge\rm \frac{ x^2 }{ 81 } + \frac{ y^2 }{ 225 } = 1\] center = ?

Nnesha (nnesha):

it's easy :-) just look at answer choices

Nnesha (nnesha):

it's horizontal or vertical ??

OpenStudy (anonymous):

c?

Nnesha (nnesha):

and how did you get that ? :-)